OpenStudy (anonymous):

Help please. Model the function rule y = -2x - 3 with a table of values.

OpenStudy (anonymous):

I have no idea how to work with functions!

jimthompson5910 (jim_thompson5910):

what happens when x = 0

jimthompson5910 (jim_thompson5910):

ie, what is the value of y when x = 0 ?

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

y = -2x - 3 y = -2(0) - 3 ... replace x with 0 (since x = 0) y = 0 - 3 y = -3

jimthompson5910 (jim_thompson5910):

so y = -3 when x = 0

jimthompson5910 (jim_thompson5910):

what is the value of y when x = 1 ?

OpenStudy (anonymous):

oh okay I get it so you replace the x with 0?

jimthompson5910 (jim_thompson5910):

yep, then you evaluate

OpenStudy (anonymous):

y = 0?

jimthompson5910 (jim_thompson5910):

no, replace x with 1 and evaluate to get y = ??

OpenStudy (anonymous):

nvm I get it! thanks(:

jimthompson5910 (jim_thompson5910):

ok tell me what you got

OpenStudy (anonymous):

if I replace x with 1 y = -5?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

just checking to make sure you got it

OpenStudy (anonymous):

thanks for checking now im 100% sure I got it

jimthompson5910 (jim_thompson5910):

sure thing